# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v2

@@ 6. Discussion
Effectiveness of Randomized Algorithms. To understand why randomized algorithms are effective in mitigating unfairness, we compare them to brute-force strategies. We create a small DNN for the Adult (Sex) dataset with 2 hidden layers, each with 16 neurons, to allow the brute force search to explore the entire space. We use a restricted search space with $n_{l}=4$ and $n_{u}=9$ that has a total of $43,076,484$ states. The time-out for SA runs is one hour. The brute force exhaustive search takes 60 hours. Table 7 highlights the results of both search strategies. SA finds the global optimal solution for two seeds in the search space. The EOD Delta column highlights the difference in cost of the states identified by SA and brute force. The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states. SA is faster than brute force in reaching a good state as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost. Additionally, SA uses a probabilistic transition function that can transition to worse states and potentially escape local minima. Gradient-based methods often get stuck in local minima. Table 8 shows the distribution of the Best, Good, and Bad states. The best states are the states with the global optimal cost. The cost of the good states is within 0.05 of the optimal cost. Bad states are states whose F1 scores are less than the 98% of the baseline F1 score. We observe that the best and good states are a negligible amount of the search space. On the other hand, over 90% of the states are bad states. This suggests that the good states occupy small “pockets” of the search space. The cost gradient effectively leads SA to identify such pockets and the probabilistic transitions prevent it from getting stuck in local minima.
Limitations. Since we used randomized algorithms, our results might be sub-optimal. While simulated annealing provides statistical guarantees on the confidence and running time to reach the optimal state, in practice such guarantees might require multiple runs of algorithms over a long period of time. Additionally, we observe a discrepancy between the validation and test fairness improvements where the test set improves less than the validation set in both NeuFair algorithms. For our experiments, we use Equalized Odds Difference (EOD) only. However, the cost function can be modified accordingly with any fairness metric.
@@ 7. Related Work
The aforementioned works tamper with the precious training data samples that might lead to unrealistic representations of different protected groups. Additionally, it may not always be feasible to modify training data. Our work is tangential as it focuses on post-processing mitigation that does not modify or access the training dataset and works on pre-trained unfair DNNs.
@@
B) In-processing techniques aim to improve fairness during training. Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions. Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the loss function to balance accuracy and fairness. These strategies necessitate alterations in either the loss function or the model parameters. Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training. Fairway (Chakraborty et al., 2020) combines pre-processing and in-processing mitigation techniques to improve fairness. Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy. While mitigating unfairness during training might be less intrusive than changing the training datasets, it often requires changing the training algorithms. This might only be effective for a particular algorithm and fairness definition. Additionally, the training process is often expensive and might not be feasible for existing systems.
@@
C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022). Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post hoc correction mechanism that modifies the output of a biased binary classifier. EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor. NeuFair, however, takes a fundamentally different approach to post-processing mitigation. First, NeuFair does not require training auxiliary models to mitigate unfairness. Second, NeuFair does not rely on randomization for fairness and generates a deterministic DNN classifier. Predictions during inference time are deterministic after the neurons are dropped from the DNN. Lastly, Woodworth et al. (2017) shows that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data. Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness. Care (Sun et al., 2022) presented a Particle Swarm Optimization (PSO) algorithm to repair DNNs. Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons. Inspired by program repair techniques, Faire first groups neurons representing protected and non-protected features in the dataset into two categories. To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute. By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected. Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted. NeuFair, however, does not retrain the model and does not separately use the protected attribute to identify unfair neurons. We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset. Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons. FairNeuron (Gao et al., 2022) proposes an algorithm to mitigate unfairness by selectively retraining neurons in a DNN. On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction. Then, the dataset is split into biased and unbiased subsets corresponding to the neuron activations in the paths. Finally, the algorithm retrains the model with both training data splits by enabling random dropouts in the biased paths and disabling any dropouts in the unbiased paths. While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair. First, we do not require access to the large training dataset, and we do not retrain the models. Second, NeuFair algorithms utilize deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly. Lastly, NeuFair does not retrain the DNN to update weights.
@@ 8. Conclusion
For future work, there are a few exciting directions. First, we can analyze the top $k$ states found during the run instead of choosing the best state. Such a strategy could reduce the difference in improvement observed between the validation and test sets. We can extend to multi-valued protected attributes (age groups) or optimize for more than one protected attribute at a time. Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models. Lastly, alternate strategies like genetic algorithms can be used to solve the combinatorial optimization problem of neuron dropout. We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems. The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution. This approach can be promising as it enables us to explore more states of the search space simultaneously.
@@
Our open-source tool NeuFair with all experimental subjects are publicly accessible: https://github.com/vdasu/neufair.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230061.

### v3

@@ 6. Discussion
Effectiveness of Randomized Algorithms. To understand why randomized algorithms are effective in mitigating unfairness, we compare them to brute-force strategies. We create a small DNN for the Adult (Sex) dataset with 2 hidden layers, each with 16 neurons, to allow the brute force search to explore the entire space. We use a restricted search space with $n_{l}=4$ and $n_{u}=9$ that has a total of $43,076,484$ states. The time-out for SA runs is one hour. The brute force exhaustive search takes 60 hours. Table 7 highlights the results of both search strategies. SA finds the global optimal solution for two seeds in the search space. The EOD Delta column highlights the difference in cost of the states identified by SA and brute force. The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states. SA is faster than brute force in reaching a good state as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost. Additionally, SA uses a probabilistic transition function that can transition to worse states and potentially escape local minima. Gradient-based methods often get stuck in local minima. Table 8 shows the distribution of the Best, Good, and Bad states. The best states are the states with the global optimal cost. The cost of the good states is within 0.05 of the optimal cost. Bad states are states whose F1 scores are less than the 98% of the baseline F1 score. We observe that the best and good states occupy a negligible amount of the search space. On the other hand, over 90% of the states are bad states. This suggests that the good states occupy small “pockets” of the search space. The cost gradient effectively leads SA to identify such pockets and the probabilistic transitions prevent it from getting stuck in local minima.
Limitations. Since we used randomized algorithms, our results might be sub-optimal. While SA provides statistical guarantees on the confidence and running time to reach an optimal state, in practice such guarantees might require multiple runs of algorithms over a long period of time. Additionally, we observe a discrepancy between the validation and test fairness improvements where the test set improves less than the validation set in both NeuFair algorithms. For our experiments, we use Equalized Odds Difference (EOD) only. However, the cost function can be modified accordingly with any fairness metric.
@@ 7. Related Work
The works above tamper with the precious training data samples that might lead to unrealistic representations of different protected groups. Additionally, it may not always be feasible to modify training data. Our work is tangential, focusing on post-processing mitigation that does not modify or access the training dataset and works on pre-trained unfair DNNs.
@@
B) In-processing techniques improve fairness during training. Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that hinders the ability of adversaries to determine sensitive attributes from ML predictions. Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the loss function to balance accuracy and fairness. These strategies necessitate alterations in either the loss function or the model parameters. Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training. Fairway (Chakraborty et al., 2020) combines pre-processing and in-processing mitigation techniques to improve fairness. Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy. While mitigating unfairness during training might be less intrusive than changing the training datasets, it often requires changing the training algorithms. This might only be effective for a particular algorithm and fairness definition. Additionally, the training process is often expensive and might not be feasible for existing systems.
@@
C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022). Equalized Odds Processing (EOP) (Hardt et al., 2016) modifies the output of a biased binary classifier to improve fairness. EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor. NeuFair, however, does not require training auxiliary models to mitigate unfairness. Additionally, NeuFair does not rely on randomization for fairness. Predictions during inference time are deterministic after the neurons are dropped from the DNN. Woodworth et al. (2017) also show that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data. Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness. Care (Sun et al., 2022) presented a Particle Swarm Optimization (PSO) algorithm to repair DNNs. Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons. Inspired by program repair techniques, Faire first groups neurons representing protected and non-protected features in the dataset into two categories. To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute. By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected features. Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted. NeuFair, however, does not retrain the model and does not separately use the protected attribute to identify unfair neurons. We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset. Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons. FairNeuron (Gao et al., 2022) proposes an algorithm to mitigate unfairness by selectively retraining neurons in a DNN. On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction. Then, the dataset is split into biased and unbiased subsets corresponding to the neuron activations in the paths. Finally, the algorithm retrains the model with both training data splits by enabling random dropouts in the biased paths and disabling any dropouts in the unbiased paths. While FairNeuron uses dropout to improve fairness, it fundamentally differs from NeuFair. We do not retrain the DNN by using the training dataset. Also, NeuFair utilizes deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly.
@@ 8. Conclusion
For future work, there are a few exciting directions. First, we can analyze the top $k$ states found during the run instead of choosing the best state. Such a strategy could reduce the difference in improvement observed between the validation and test sets. We can extend to multi-valued protected attributes (e.g., age groups) or optimize for more than one protected attribute at a time (e.g., race and sex). Second, we plan to leverage dropouts in the fine-tuning process to understand the effect of randomized algorithms on mitigating the unfairness and toxicity of large language models. Lastly, alternate strategies like genetic algorithms may be used to solve the combinatorial optimization problem of neuron dropout. We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems. The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution. This approach can be promising as it enables us to explore more states of the search space simultaneously.
@@
Our open-source tool NeuFair with all experimental subjects are publicly accessible on Zenodo (Dasu, 2024) and GitHub.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230060 and CNS-2230061.